OpenStudy (raquelkelli01):

Indicate the equation of the line through (2, -4) and having slope of 3/5.

OpenStudy (alexandervonhumboldt2):

y=mx+b form where m is slope: m=3/5 substitute 3/5 for slope: y=3/5x+b now substitute the point and solve for b: -4=3/5*2+b b=?

OpenStudy (mathmale):

This is a perfectly acceptable way to find the desired equation. Another would be the "point-slope" formula, \[y-y _{1}=m(x-x _{1}).\]

OpenStudy (mathmale):

Please note that "y=3/5x+b" could be written better as y=(3/5)x+b. Less ambiguous that way.